### v2

# Quantum simulation of many-body dynamics with noise-robust Trotter decomposition based on symmetric structures

Abstract
The Suzuki-Trotter decomposition, which digitalizes quantum time evolution, provides a promising framework for simulating quantum dynamics on quantum hardware and exploring quantum advantage over classical computation. However, conventional Trotter circuits require a large number of non-local gates, lowering their faithfulness to the ideal dynamics when implemented on current noisy quantum hardware. While most previous studies have focused on circuit optimization, we instead propose a new Trotter decomposition that is intrinsically circuit-efficient for simulating quantum dynamics on near-term devices. Our method substantially reduces the number of CNOT operations compared to conventional Trotter decompositions by exploiting the symmetry of the target model to construct an effective Hamiltonian with fewer two-qubit gates. We demonstrate the noise robustness of the proposed approach through numerical simulations of a nine-site Heisenberg model under realistic noise, and further validate its experimental practicality on the IBM superconducting device, achieving a state fidelity exceeding $0.98$ when combined with quantum error mitigation in the three-site case. The proposed circuit design is also compatible with existing circuit optimization techniques. Our results establish a practical route toward noise-resilient quantum simulation in many-body dynamics.
preprint: APS/123-QED
I Introduction
Through the rapid advance of quantum hardware, quantum simulation has gained emerging attention to simulate physical and chemical models that become practically intractable by classical computational resources Lanes et al. (2025); Huang et al. (2025). Its prominent and versatile applicability lies in non-equilibrium quantum many-body dynamics, where the digital simulation based on Suzuki-Trotter decomposition enables tractable and scalable approximation of real-time evolution Trotter (1959); Suzuki (1976); Hatano and Suzuki (2005). Simulating Trotter steps with quantum computers frees from classical simulation with exponential computational resources, requiring only linear overhead to system size, which is thus seen as one of the applications with potential near-term quantum advantage.
However, the non-negligible noise level and hardware restrictions of current quantum hardware still pose a significant obstacle to the practical realization of such Trotter-based simulations. In particular, superconducting quantum devices ibm (2022), which are among the most extensively developed and commercially accessible platforms, suffer from noisy non-local gates and limited coherence times Mooney et al. (2021); Yang et al. (2022a). Therefore, it is essential to design quantum circuits with reduced depth and fewer non-local gates, such as CNOT gates, to alleviate noise accumulation and improve the fidelity of simulations.
While substantial efforts have focused on optimizing given Trotter circuits under hardware constraints Lötstedt et al. (2023); Chowdhury et al. (2024a); Choi, Chowdhury, and Yu (2025); Yang et al. (2022b); Zhao et al. (2023); Jones et al. (2019), the underlying Trotter decomposition itself sets the fundamental limits of such optimization. In this work, we design a new alternative Trotter decomposition strategy that substantially reduces the number of CNOT gates in use. We exploit the symmetric structure of the given Hamiltonian, particularly, of the $XXX$ Heisenberg model. We transform the three-site Heisenberg Hamiltonian into a more concise two-site effective Hamiltonian through an encoding and decoding procedure, thereby enabling more efficient circuit construction.
This new decomposition with the effective Hamiltonian reduces the average number of CNOT gates in each Trotter step to $2.625$ per qubit, whereas the conventional Trotter circuit requires $3$. The proposed method highlights the potential of reducing the circuit overhead with a more efficient approach for Trotter decomposition rather than merely performing circuit optimization on existing Trotter circuits.
We demonstrate the noise-robustness of our proposed method through numerical simulation. Our method outperforms the conventional Trotter decomposition in simulating the time evolution of the nine-site $XXX$ Heisenberg model under depolarizing noise. We also simulate the time evolution of a three-site $XXX$ Heisenberg model on the superconducting quantum device ibmq_jakarta provided by the IBM Quantum Platform ibm (2022). Using quantum error mitigation (QEM) Li and Benjamin (2017); Temme, Bravyi, and Gambetta (2017); Endo, Benjamin, and Li (2018); Giurgica-Tiron et al. (2020); Yang, Raymond, and Uno (2022); Yang et al. (2023); Cai et al. (2023); Endo et al. (2021), we achieve the target state fidelity over $0.98$ on ibmq_jakarta. In implementing the proposed Trotter circuits, our method finds further compatibility with circuit optimization with the Qiskit package Javadi-Abhari et al. (2024). The overall experiments demonstrate that our method offers not only a novel, efficient Trotter decomposition scheme but also a practical and feasible solution for simulating physical models on current quantum hardware.
II Suzuki-Trotter Decomposition
We consider $N$-site $J=1$ XXXXXX Heisenberg Hamiltonian with $N=4M+1$, $M\in\mathbb{Z}_{\geq 0}$, and open boundary condition formalized as
$\hat{H}=\sum_{i=1}^{N-1}\vec{\sigma}^{(i)}\cdot\vec{\sigma}^{(i+1)},$ (1)
where $\cdot$ denotes the inner product of three components of the Pauli operators of $i$-th site $\{\hat{\sigma}_{x}^{(i)},\hat{\sigma}_{y}^{(i)},\hat{\sigma}_{z}^{(i)}\}$ defined as
$\vec{\sigma}^{(i)}\cdot\vec{\sigma}^{(j)}:=\sum_{\mu\in\{x,y,z\}}\hat{\sigma}_{\mu}^{(i)}\otimes\hat{\sigma}^{(j)}_{\mu}\quad\text{for}\quad i\not=j.$ (2)
Given a Heisenberg Hamiltonian $\hat{H}=\hat{O}_{1}+\hat{O}_{2}$ that consists of two non-commutative operators $\hat{O}_{1}$ and $\hat{O}_{2}$, i.e. $\hat{O}_{1},\hat{O}_{2}]\not=0$, the conventional Trotter decomposition with $n$ steps Trotter (1959); Suzuki (1976); Hatano and Suzuki (2005) approximates the evolution of this Hamiltonian with
$\hat{U}(t)=\left(\exp(-{\mathrm{i}}\hat{O}_{2}\Delta t)\exp(-{\mathrm{i}}\hat{O}_{1}\Delta t)\right)^{n}+\mathcal{O}(Mn^{-1}),$ (3)
where $\Delta t=t/n$ denotes the evolution time for a step, and $\hat{O}_{1}$ and $\hat{O}_{2}$ are chosen to be
$\hat{O}_{1}$ $=\sum_{i=1}^{2M}\vec{\sigma}^{(2i-1)}\cdot\vec{\sigma}^{(2i)},$ (4)
$\hat{O}_{2}$ $=\sum_{i=1}^{2M}\vec{\sigma}^{(2i)}\cdot\vec{\sigma}^{(2i+1)}.$ (5)
The quantum circuit for the decomposition Eq. (3) can be constructed in the following way. First, let $\hat{u}^{(i)}(\Delta t)$ be the unitary operator of evolution time $\Delta t$ for $i$-th and $i+1$-th qubit defined as
$\hat{u}^{(i)}(\Delta t)=\exp(-{\mathrm{i}}\vec{\sigma}^{(i)}\cdot\vec{\sigma}^{(i+1)}\Delta t).$ (6)
Since any two-qubit unitary operation can be realized by three CNOT gates Vatan and Williams (2004); Vidal and Dawson (2004); Shende, Bullock, and Markov (2006), Fig. 1 provides a quantum circuit to implement $\hat{u}^{(i)}(t)$.
Figure 1: The quantum circuit of the unitary operator in Eq. (6).
Using this circuit block, the Trotter decomposition in Eq. (3) is then constructed by the quantum circuit in Fig. 2. Since each time step has two layers of operator $\hat{u}^{(i)}(t)$, the averaged number of CNOT gates applied over each qubit is three, which represents the most efficient circuit construction regarding the CNOT overhead known to date Chowdhury et al. (2024b).
Figure 2: The quantum circuit of the time evolution operator using the Trotter decomposition in Eq. (3) for $N=5$. The operation corresponding to the propagation of each single time step $\Delta t$ is enclosed by a dashed box.
III Proposed Decomposition
In order to reduce the number of CNOT gates in each single time step, we propose a new Trotter decomposition. First, we change the way of partitioning the Heisenberg Hamiltonian into $\hat{H}=\hat{O^{\prime}}_{0}+\hat{O^{\prime}}_{1}+\hat{O^{\prime}}_{2}$ with the operators
$\hat{O}^{\prime}_{0}$ $=\sum_{i=1}^{M}\vec{\sigma}^{(4i-2)}\cdot\vec{\sigma}^{(4i-1)}+\vec{\sigma}^{(4i-1)}\cdot\vec{\sigma}^{(4i)},$ (7)
$\hat{O}^{\prime}_{1}$ $=\sum_{i=1}^{M}\vec{\sigma}^{(4i-3)}\cdot\vec{\sigma}^{(4i-2)},$ (8)
$\hat{O}^{\prime}_{2}$ $=\sum_{i=1}^{M}\vec{\sigma}^{(4i)}\cdot\vec{\sigma}^{(4i+1)}.$ (9)
Using $\hat{O}^{\prime}_{0}$, $\hat{O}^{\prime}_{1}$, and $\hat{O}^{\prime}_{2}$ with $m$ proposed Trotter blocks for evolution time $\Delta t^{\prime}=t/m$ is then defined as
$\begin{split}&\hat{U}_{\mathrm{Heis}}(t)\\ &=\left(\exp(-{\mathrm{i}}\hat{O}^{\prime}_{1}\Delta t^{\prime})\exp(-{\mathrm{i}}\hat{O}^{\prime}_{0}\Delta t^{\prime})\exp(-{\mathrm{i}}\hat{O}^{\prime}_{2}\Delta t^{\prime})\right)^{m}\\ &\quad+\mathcal{O}\left(Mm^{-1}\right).\end{split}$ (10)
Remarkably, the number of required time steps $m$ in Eq. (10) is smaller than $n$ in Eq. (3) for the same residual error of Trotter decomposition. This arises from the fact that Eq. (10) implements the time evolution operator $e^{-{\mathrm{i}}\hat{O^{\prime}}_{0}\Delta t^{\prime}}$ directly instead of decomposing it into $\prod_{i=1}^{M}e^{-{\mathrm{i}}\vec{\sigma}^{(4i-2)}\cdot\vec{\sigma}^{(4i-1)}\Delta t^{\prime}}e^{-{\mathrm{i}}\vec{\sigma}^{(4i-1)}\cdot\vec{\sigma}^{(4i)}\Delta t^{\prime}}$ that causes additional error in the order of $\mathcal{O}(Mm^{-1})$. This implies that the residual error of Eq. (10) is reduced to three-quarters of that of the conventional Trotter decomposition for $m=n$. As a result, using $m=\frac{3}{4}n$ steps of the proposed Trotter blocks achieves the same level of residual error as the conventional Trotter blocks.
In designing a quantum circuit for this new Trotter decomposition, we can use $\{\hat{u}^{(4i-1)}(\Delta t^{\prime})\}_{i=1}^{M}$ and $\{\hat{u}^{(4i)}(\Delta t^{\prime})\}_{i=1}^{M}$ to construct the time evolution operators $\exp(-{\mathrm{i}}\hat{O}_{1}^{\prime}\Delta t^{\prime})$ and $\exp(-{\mathrm{i}}\hat{O}_{2}^{\prime}\Delta t^{\prime})$ in Eq. (10) since they share the same form as Eq. (6). Thus, it suffices to design an efficient circuit of the time evolution operator $\exp(-{\mathrm{i}}\hat{O}^{\prime}_{0}\Delta t^{\prime})$, where $\hat{O}_{0}^{\prime}$ takes the form of a three-site $XXX$ Heisenberg Hamiltonian $\hat{H}_{3}=\vec{\sigma}^{(1)}\cdot\vec{\sigma}^{(2)}+\vec{\sigma}^{(2)}\cdot\vec{\sigma}^{(3)}$.
To construct an efficient quantum circuit of $\exp(-{\mathrm{i}}\hat{H}_{3}t)$, we aim to compress it to a smaller subsystem by deriving an effective Hamiltonian $H_{\mathrm{eff}}$. We focus on the fact that $\hat{H}_{3}$ commutes with $\hat{\sigma}_{\mu}^{(1)}\otimes\hat{\sigma}_{\mu}^{(2)}\otimes\hat{\sigma}_{\mu}^{(3)}$, i.e. $\hat{H}_{3},\hat{\sigma}_{\mu}^{(1)}\otimes\hat{\sigma}_{\mu}^{(2)}\otimes\hat{\sigma}_{\mu}^{(3)}]=0$. This yields simultaneous eigenstates as follows:
[…]
Figure 3: The quantum circuit to realize the encoder $\hat{U}_{\mathrm{enc}}$.
The encoder $\hat{U}_{\mathrm{enc}}$ then transforms $\hat{H}_{3}$ to a two-qubit effective Hamiltonian $\hat{H}_{\mathrm{eff}}$,
$\begin{split}\hat{H}_{\mathrm{eff}}&=\hat{U}_{\mathrm{enc}}\hat{H}_{3}\hat{U}_{\mathrm{enc}}^{\dagger}\\ &=\hat{\sigma}_{x}^{(1)}+\hat{\sigma}_{x}^{(2)}+\hat{\sigma}_{z}^{(1)}+\hat{\sigma}_{z}^{(2)}-(\hat{\sigma}_{z}^{(1)}\otimes\hat{\sigma}_{x}^{(2)}+\hat{\sigma}_{x}^{(1)}\otimes\hat{\sigma}_{z}^{(2)}).\end{split}$ (15)
This yields the following equivalence between the time evolution operators:
$\begin{split}\exp\left(-{\mathrm{i}}\hat{H}_{3}\Delta t^{\prime}\right)&=\hat{U}_{\mathrm{enc}}^{\dagger}\exp\left(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}\right)\hat{U}_{\mathrm{enc}}.\end{split}$ (16)
This suggests that the time evolution under $\hat{H}_{3}$ can be realized as the composition of the encoder $\hat{U}_{\mathrm{enc}}$, the time evolution under $\hat{H}_{\mathrm{eff}}$, and the decoding unitary $\hat{U}_{\mathrm{enc}}^{\dagger}$.
Figure 4: The quantum circuit of the time evolution operator $e^{-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}}$ given by Eq. (17).
To design an efficient circuit implementation of $\exp(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime})$, we further expand $\exp(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime})$ in the following form
$\begin{split}&\exp(-i\hat{H}_{\mathrm{eff}}\Delta t^{\prime})\\ &=e^{-{\mathrm{i}}(\hat{\sigma}_{z}^{(1)}+\hat{\sigma}_{x}^{(2)})\Delta t^{\prime}}e^{{\mathrm{i}}(\hat{\sigma}_{z}^{(1)}\otimes\hat{\sigma}_{x}^{(2)}+\hat{\sigma}_{z}^{(2)}\otimes\hat{\sigma}_{x}^{(1)})\Delta t^{\prime}}e^{-{\mathrm{i}}(\hat{\sigma}_{z}^{(2)}+\hat{\sigma}_{x}^{(1)})\Delta t^{\prime}}\\ &\quad+\mathcal{O}(m^{-3}),\end{split}$ (17)
where the residual error decreases faster than that of the conventional Trotter decomposition. This provides a quantum circuit of the time evolution $\exp(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime})$, represented in Fig. 4.
Figure 5: (a) A concise form of the quantum circuit of the time evolution operator, denoting $\Delta t^{\prime}$ time evolution. (b) The quantum circuit of the new notation of the two-qubit unitary gate $\hat{u}^{\prime}$ seen in (a).
Using the circuit implementation of $\exp(-{\mathrm{i}}\hat{O}^{\prime}_{0}\Delta t^{\prime})$, the whole quantum circuit of the proposed Trotter block for $\Delta t^{\prime}$ in Eq. (10) becomes Fig. 5(a), which contains $14$ CNOT gates in each four-qubit Trotter block. Thus, our circuit construction consumes on average $3.5$ CNOT gates per qubit in a single Trotter block with evolution time $\Delta t^{\prime}$. Moreover, taking the ratio between $\Delta t$ and $\Delta t^{\prime}$ into account, our circuit construction requires only $3.5m=2.625n$ CNOT gates to perform $t=n\Delta t$ time evolution for. Since the conventional Trotter circuit requires an average of $3n$ CNOT gates per qubit up to time $t$, yielding a reduction rate of $2.625n/3n=0.875$ compared to the original Trotter blocks. This reduction of CNOT gates significantly contributes to the noise resilience of our proposed Trotter decomposition.
Figure 6: Comparison between the conventional Trotter decomposition (Fig. 1) and the proposed decomposition (Fig. 5), where plots of the conventional Trotter decomposition are colored red and those of the proposed decomposition are colored blue. (a) The state infidelity of the evolved state under the depolarizing noise with $p_{1}=1.0\times 10^{-4}$ to the ideally evolved state without noise, scaling with different numbers of Trotter steps $n$. Note that the x-axis for the proposed method is normalized to $3m/4$ on the plot so that the residual error of Eq. (3) and Eq. (10) become compatible. (b) The state infidelity of the evolved state under the noise to the ideally evolved state without noise, scaling with different depolarizing probabilities $p_{1}\in\{1.0\times 10^{-5},3.0\times 10^{-4},1.0\times 10^{-4},3.0\times 10^{-4}\}$ and $p_{2}=10p_{1}$.
IV Experiments
To demonstrate the practicality of the proposed new Trotter decomposition, we simulate the time evolution of the $XXX$ Heisenberg model and calculate the fidelity $F(\tilde{\rho},\rho_{\mathrm{ideal}})=\operatorname{Tr}\left[\left(\rho_{\mathrm{ideal}}^{1/2}\tilde{\rho}\rho_{\mathrm{ideal}}^{1/2}\right)^{1/2}\right]^{2}$ of the resulting state $\tilde{\rho}$ to the ideally evolved state $\rho_{\mathrm{ideal}}$. We use the noisy simulator, real-device emulator fake_jakarta, and real quantum device ibmq_jakarta provided by IBM Quantum Platform.
IV.1 Numerical simulation under depolarizing noise
Figure 7: (a) The number of the CNOT gates scaling with different numbers of Trotter steps $n$. Note that the x-axis for the proposed method is normalized to $3m/4$ on the plot so that the residual error of Eq. (3) and Eq. (10) become compatible. (b) The infidelity between the evolved state under the noise and the ideally evolved state without noise, scaling with different Trotter steps.
Figure 8: (a) The infidelity of the evolved state to the expected noise-free state, simulated by the two encoding-decoding strategies among different Trotter steps $m$. (b) The infidelity of the evolved state to the expected noise-free state, for different QEM levels among different Trotter steps $m$.
We first perform the noisy numerical simulation of the nine-site $XXX$ Heisenberg model up to a fixed evolution time $t=\pi$. We introduce depolarizing noise in both single-qubit gates and two-qubit gates with depolarizing probabilities $p_{1}=1.0\times 10^{-4}$ and $p_{2}=10p_{1}=1.0\times 10^{-3}$, respectively, which reflect the typical noise levels of the current quantum hardware. Starting from an arbitrary initial state, which we set to $|110100110\rangle$, we compare the state fidelity between the conventional Trotter circuit in Fig. 1 and the proposed Trotter circuit in Fig. 5 with different Trotter steps among $n,m\in\{4,8,12,\ldots,96\}$.
The simulated results between the conventional Trotter decomposition and the proposed decomposition are plotted in Fig. 6. We observe from Fig. 6(a) that the proposed method achieves lower state infidelity than the conventional approach. We also observe that there exists an optimal number of Trotter steps that balances the simulation accuracy and the noise effect induced when increasing the Trotter steps. Focusing on the infidelity with the optimal number of Trotter steps, Fig. 6(b) visualizes the advantage of our proposed method over the conventional method in terms of lower infidelity among different noise levels.
The results mentioned above imply that our proposed decomposition is more noise-robust thanks to its reduced use of CNOT gates. Fig. 7(a) visualizes the number of CNOT gates used in the whole quantum circuit between the proposed method and the conventional one, reflecting the estimated reduction ratio $0.875$. We further plot the relation between the number of CNOT gates and the state infidelity in Fig. 7(b), where the proposed method achieves smaller state infidelities with the same number of CNOT gates. This implies that the proposed quantum circuit constructs more noise-resilient Trotter blocks than the conventional circuit.
IV.2 Real-device experiments with error suppression
Figure 9: The error map of ibmq_jakarta on April 16, 2022. The numbers on the figure represent the indices of physical qubits. We use the physical qubits 5, 3, and 1 with the virtual qubit indices 0, 1, and 2 on quantum circuits. The device noise is subject to temporal fluctuations.
[…]
Based on the above, we compute the infidelity of the resulting state at the evolution time $t=\pi$ evolved from a given initial state $|110\rangle$ at $t=0$. We add quantum error mitigation (QEM) Li and Benjamin (2017); Temme, Bravyi, and Gambetta (2017); Endo, Benjamin, and Li (2018); Giurgica-Tiron et al. (2020); Yang, Raymond, and Uno (2022); Yang et al. (2023); Cai et al. (2023); Endo et al. (2021) to reduce the noise effect through classical post-processing. Particularly, we use quantum readout error mitigation (QREM) Yang, Raymond, and Uno (2022) and zero-noise extrapolation (ZNE) Temme, Bravyi, and Gambetta (2017); Giurgica-Tiron et al. (2020). We use the digital ZNE method Giurgica-Tiron et al. (2020) with the linear fitting method and the scale factors $1.0$, $2.0$ and $3.0$, provided by Mitiq LaRose et al. (2022). Furthermore, the Pauli twirling technique Bennett et al. (1996); Knill (2004); Wallman and Emerson (2016) is also combined with ZNE, referring to the implementation by Berthusen et al. Berthusen et al. (2021). Each quantum circuit is executed with $8192$ shots, and the infidelity is averaged over $8$ samples.
First, we examine the performance among Trotter steps $\{4,5,6,7,8,9,10,20,30,40\}$ on the noisy simulator fake_jakarta, comparing the two encoding-decoding methodologies: the general encoder $\hat{U}_{\mathrm{enc}}$ and general decoder $\hat{U}_{\mathrm{enc}}^{\dagger}$ (general-general), and the aforementioned shallow encoding and specific encoding (shallow-specific). Here, we apply only QREM to noisy results before computing the infidelity. The result is shown in Fig. 8(a), where both encoding-decoding methods achieve the infidelity below $0.2$ with more than $10$ Trotter steps and the shallow-specific method further achieves the infidelity smaller than $0.1$.
The effect of QEM methods is also investigated under fake_jakarta. Here, we set the configuration to shallow encoding and specific decoding. We observe from Fig. 8(b) that both QREM and ZNE contribute to reducing the infidelity. The instability of ZNE can be improved by adding Pauli twirling that tailors the noise to a stochastic Pauli channel. We also see that combining Pauli twirling further enhances the accuracy gains achieved by finer Trotter step decomposition.
Settings fake_jakarta ibmq_jakarta
General-general
Without QEM $0.7856\pm 0.0015$ $0.8039\pm 0.0048$
QREM $0.8448\pm 0.0015$ $0.9032\pm 0.0054$
QREM, ZNE $0.9393\pm 0.0053$ $0.9866\pm 0.0017$
QREM, ZNE, Twirling $0.9801\pm 0.0031$ -
Shallow-specific
Without QEM $0.8631\pm 0.0017$ $0.8637\pm 0.0041$
QREM $0.9234\pm 0.0016$ $0.9728\pm 0.0040$
QREM, ZNE $0.9840\pm 0.0024$ $0.9857\pm 0.0043$
QREM, ZNE, Twirling $0.9714\pm 0.0048$ $0.9624\pm 0.0167$
Table 1: The fidelity of the simulated state from IBM Quantum Jakarta and its fake simulator under different QREM levels and encoding-decoding strategies. “General-general” represents the use of the general encoder and the general decoder, and “Shallow-specific” represents the use of the shallow encoder and the specific decoder.
Finally, we examine the time evolution from $t=0$ to $t=\pi$ on the real quantum device ibmq_jakarta. We execute $100$ Trotter steps with and without QEM under the two encoding-decoding methods. The state fidelity is then calculated by reconstructing the density matrix through state tomography Raymer, Beck, and McAlister (1994); Leonhardt (1995); Leibfried et al. (1996).
The results are listed in Table 1, where the fidelities obtained by fake_jakarta and ibmq_jakarta are compared. All the fidelities exceed $0.80$ on ibmq_jakarta, and all the fidelities even exceed $0.90$ with QREM only. Remarkably, we achieve a fidelity over $0.98$ with ZNE with the general-general method, which ensures the generality of our method in simulating the dynamics from an arbitrary initial state. All the experimental results on the noisy simulator and the real device support the practicality of our proposed Trotter decomposition.
V Conclusion
In this work, we propose a novel, noise-resilient Trotter decomposition focusing on the symmetry of the given Heisenberg Hamiltonian to notably reduce the number of CNOT gates in its circuit implementation without sacrificing the accuracy of the Trotter decomposition. The noise-robustness of the proposed method is demonstrated through the numerical simulation under noise. Our experiments also record high fidelities in simulating the three-site Heisenberg model on the real quantum device. The proposed method thus establishes a direct link between physical insight into the model’s symmetry and quantum circuit design regarding efficient Trotter decomposition, offering both a fresh theoretical perspective and practical benefits for performing noise-resilient quantum dynamics simulation.
This work opens up several promising directions. First, thanks to the generality of the proposed approach, it can be extended to a broader range of physical and chemical models, such as transverse-field Ising models and Fermi-Hubbard models. Next, integrating more recent QEM techniques based on ZNE Endo et al. (2019); Hakkaku et al. (2025) would further suppress both coherent and algorithmic errors in Trotterized circuits. Moreover, resource-efficient implementations using subspace expansion methods Yoshioka et al. (2022); Yang et al. (2023) can also be adapted to our framework to mitigate both coherent and stochastic errors. In combination with quantum–classical divide-and-conquer approaches Sun et al. (2022); Eddins et al. (2022); Yuan et al. (2021); Harada et al. (2025), these techniques would further enhance the practical feasibility of our method on near-term quantum devices.
Acknowledgements.
The result of real-device experiments is obtained as a solution to the IBM Quantum Awards: Open Science Prize 2021, for which the initial manuscript has been publicly available at Yang and Negishi (2022) since April 2022. B.Y. and N.N. sincerely thank all those who made this contest possible.
Data Availability Statement
The code and data used in the experiments in this work are publicly available on the GitHub page: https://github.com/BOBO1997/osp_solutions Yang and Negishi (2022).